A certain kind of sheet metal has, on average, 9 defects per 14 square feet.

Assuming a Poisson distribution, find the probability that a 25 square foot metal sheet has at least 14 defects. Round your answer to four decimals.


1
Expert's answer
2022-03-02T12:48:10-0500

Let "X=" the number of defects: "X\sim Po(\lambda)."


"P(X=k)=\dfrac{e^{-\lambda}(\lambda)^k}{k!}"

Given "\lambda=\dfrac{25}{14}\cdot 9\approx16.07143"


"P(X\ge14)=1-P(X<14)"

"=1-P(X=0)-P(X=1)-P(X=2)"

"-P(X=3)-P(X=4)-P(X=5)"

"-P(X=6)-P(X=7)-P(X=8)"

"-P(X=9)-P(X=10)-P(X=11)"

"-P(X=12)-P(X=13)\approx0.3609"